OpenStudy (kkutie7):

The Mean Value Theorem for Integrals states that for a continuous and positive function f on a closed interval [a,b], there must exist a number c in [a,b], such that: Find all values of c in the interval [0,90], which satisfies the Mean Value Theorem for Integrals for the function R(t) given in Question 9. Show the equation needed to be solved to find the value of c. (Note: You may find c using any method you choose. That is, you can solve the equation graphically and/or with a CAS.)

OpenStudy (kkutie7):

\[f(c)=\frac{ 1 }{ b-a }\int\limits_{a}^{b}f(x)dx\]

zepdrix (zepdrix):

Where is R(t) from question 9? >.<

OpenStudy (kkutie7):

@zepdrix R(t) = .4t − 20 cos(π t / 45) + 40

OpenStudy (kkutie7):

Is that really it? that seems so much more simple than what's going on in my head.

OpenStudy (anonymous):

since apparently the answer is \(58\) our job it to solve \[.4t − 20 cos(π t / 45) + 40 =58\]

OpenStudy (kkutie7):

t here would be c for f(c) though.... right?

OpenStudy (anonymous):

solve for \(t\) use wolfram

OpenStudy (kkutie7):

alright I'll try that.

OpenStudy (anonymous):

Sattty liiiite why u do dis to me my love